OpenStudy (anonymous):

Find all values of x in the interval 0,2pi that satisfy the equation 2+cos2x=3 cosx

OpenStudy (anonymous):

Well, 0 is one that satisfy that equation 2+cos(0)=3cos(0) 2+1=3*1 3=3

OpenStudy (alekos):

wish it was that easy! you need to use the identity cos2x=2cos^2x+1 and then you get a quadratic which needs to be solved for cosx

OpenStudy (anonymous):

Ohh, now it looks easy...

OpenStudy (alekos):

have you worked it out yet?